Find the total number of seats of an amphitheater if it has 70 rows with 40 seats, 44 seats and 48 seats on the first, second and third rows respectively.


1
Expert's answer
2020-10-04T18:23:13-0400

To get the total number of seats;


"=\frac{n} {2} (2a+(n-1)d)" where "n" is the total number of rows, "d" is the common difference between the rows and "a" is the number of seats in the first row.

a=40

n=70

d=4


"=\frac{70} {2} (2×40+(70-1)4)"


=38(80"+" 276)


=38(356)


=12,460 seats.